Times-Interest-Earned Ratio

The Morris Corporation has $1,000,000 of debt outstanding, and it pays an interest rate of 10% annually. Morris's annual sales are $5 million, its average tax rate is 35%, and its net profit margin on sales is 3%. If the company does not maintain a TIE ratio of at least 3 to 1, its bank will refuse to renew the loan and bankruptcy will result. What is Morris's TIE ratio? Round intermediate calculations to two decimal places. Round your answer to two decimal places.

Last year Mason Inc. had a total assets turnover of 1.33 and an equity multiplier of 1.75. Its sales were $375,000 and its net income was $10,549. The CFO believes that the company could have operated more efficiently, lowered its costs, and increased its net income by $5,250 without changing its sales, assets, or capital structure. Had it cut costs and increased its net income in this amount, by how much would the ROE have changed?

Select the correct answer.

Times-Interest-Earned Ratio

The Morris Corporation has $1,000,000 of debt outstanding, and it pays an interest rate of 10% annually. Morris's annual sales are $5 million, its average tax rate is 35%, and its net profit margin on sales is 3%. If the company does not maintain a TIE ratio of at least 3 to 1, its bank will refuse to renew the loan and bankruptcy will result. What is Morris's TIE ratio? Round intermediate calculations to two decimal places. Round your answer to two decimal places.

Last year Mason Inc. had a total assets turnover of 1.33 and an equity multiplier of 1.75. Its sales were $375,000 and its net income was $10,549. The CFO believes that the company could have operated more efficiently, lowered its costs, and increased its net income by $5,250 without changing its sales, assets, or capital structure. Had it cut costs and increased its net income in this amount, by how much would the ROE have changed?

Select the correct answer.

a. 3.26%

b. 2.08%

c. 1.49%

d. 3.85%

e. 2.67%


Future Value of a Single Payment

If you deposit $10,000 in a bank account that pays 8.1% interest annually, how much would be in your account after 5 years? Round your answer to the nearest cent.

$  

Present Value of a Single Payment

What is the present value of a security that will pay $20,000 in 20 years if securities of equal risk pay 5.3% annually? Round your answer to the nearest cent.

$  

Your parents will retire in 18 years. They currently have $240,000, and they think they will need $1 million at retirement. What annual interest rate must they earn to reach their goal, assuming they don't save any additional funds? Round your answer to two decimal places.

%

Number of Periods of a Single Payment

If you deposit money today in an account that pays 6.1% annual interest, how long will it take to double your money? Round your answer to the nearest whole number.

years

Future Value of a Single Payment

If you deposit $10,000 in a bank account that pays 8.1% interest annually, how much would be in your account after 5 years? Round your answer to the nearest cent.

$  

Present Value of a Single Payment

What is the present value of a security that will pay $20,000 in 20 years if securities of equal risk pay 5.3% annually? Round your answer to the nearest cent.

$  

Your parents will retire in 18 years. They currently have $240,000, and they think they will need $1 million at retirement. What annual interest rate must they earn to reach their goal, assuming they don't save any additional funds? Round your answer to two decimal places.

%

Number of Periods of a Single Payment

If you deposit money today in an account that pays 6.1% annual interest, how long will it take to double your money? Round your answer to the nearest whole number.

years
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Interest = Debt * Rate

= 1000000*10% = 100,000

Net Income = Net Profit margin* Sales

= 3%*5000000

=150000

Profit before tax = 150000*100/65=230769.2

EBIT = Profit before tax + Interest

= 230769.2+ 100,000

= 330769.2

TIE = EBIT/ Interest

= 330769.2/ 100000

= 3.31

2: Total assets = Sales/TAT

=375000/1.33=281954.9

Equity Multiplier= Total assets/ Equity

1.75= 281954.9/ Equity

Equity = 281954.9/1.75 = 161117.1

ROE = Net Income/ Equity = 10549/161117.1

=6.55%

New ROE = (10549+5250)/161117.1

=9.81%

Change in ROE = 9.81%-6.55% = 3.26%

3: FV = PV*(1+r)^n

= 10000*1.081^5

=14761.43

4: PV = FV/(1+r)^n

=20000/1.053^20

=7119.71